OpenStudy (anonymous):

(-5)/(x-3) + (-4)/(x+2)

OpenStudy (anonymous):

https://www.connexus.com/content/media/478541-3282011-32814-PM-1583904793.jpg what it looks like

OpenStudy (anonymous):

(-5)/(x-3) + (-4)/(x+2) -5x-10=-4x+12 +4x +4x x-10=12 +10 +10 x=22 I AM NOT SURE

OpenStudy (anonymous):

PUT the answer of x in the place of x and add